# Convergence and Generalization of Anti-Regularization for Parametric Models

### v1

@@ C.2 Proof of Test Risk Improvement Theorem
Subtract the test risks of the two models:
$\mathbb{E}\left[\mathcal{R}(\hat{f}_{\lambda})-\mathcal{R}(\hat{f}_{0})\right]=\mathbb{E}\left[\widehat{\mathcal{R}}_{S}(\hat{f}_{\lambda})-\widehat{\mathcal{R}}_{S}(\hat{f}_{0})\right]+\frac{2\tau^{2}}{n}\left(\operatorname{tr}(S_{\lambda})-\operatorname{tr}(S_{0})\right).$
The second term is $\tfrac{2\tau^{2}}{n}\Delta\operatorname{tr}$, and the first is $\Delta\widehat{\mathcal{R}}_{S}$. If $\Delta\widehat{\mathcal{R}}_{S}>\tfrac{2\tau^{2}}{n}\Delta\operatorname{tr}$, their sum is negative. ∎
@@
At equality, the two risks are the same; below that threshold, improvement occurs. The term $\tfrac{2\tau^{2}}{n}\operatorname{tr}(S_{\lambda})$ represents a variance penalty, so increasing the degrees of freedom too much via $\lambda$ (near the safe boundary) can nullify the benefit.
@@ C.3 Monotonicity and Sensitivity Analysis
so $\operatorname{tr}(S_{\lambda})$ is strictly increasing in $\lambda$.
@@
We apply termwise differentiation to $\operatorname{tr}(S_{\lambda})=\sum_{j}\tfrac{\sigma_{j}}{\sigma_{j}-\lambda}$. ∎
Corollary 20 (Lipschitz-Type Bounds).
@@
Moreover, each term satisfies
$\frac{\sigma_{j}}{(\sigma_{j}-\lambda)^{2}}\geq\frac{s_{\min}}{(s_{\max}-\lambda)^{2}},$
so
@@
The sensitivity of each term with respect to $\sigma_{j}$ is given by
$\frac{\partial}{\partial\sigma_{j}}\left(\frac{\sigma_{j}}{\sigma_{j}-\lambda}\right)=-\frac{\lambda}{(\sigma_{j}-\lambda)^{2}}<0.$
Thus, the larger the small eigenvalues $\sigma_{j}$, the more suppressed the degree-of-freedom inflation; smaller $\sigma_{j}$ lead to sharper increases in $\operatorname{tr}(S_{\lambda})$.
@@
When many small $\sigma_{j}$ exist, $\operatorname{tr}(S_{\lambda})$ becomes steeper, leading to rapid growth of the variance penalty. Combining a safe slack margin $\lambda\leq(1-\epsilon)s_{\min}$ with a sample-size-based schedule $|\lambda(n)|\propto n^{-\alpha}$ helps keep the degrees of freedom per sample under control.
Appendix D Sample Size Scaling Rule (Regression)
D.1 Degree-of-Freedom Targeting Rule
We aim to maintain a fixed complexity per sample by solving
@@ D.1 Degree-of-Freedom Targeting Rule
for $\lambda$, where $S_{\lambda}=X(X^{\top}X-n\lambda I)^{-1}X^{\top}$, and the eigenvalues of $\hat{\Sigma}=\tfrac{1}{n}X^{\top}X$ are denoted $\{\sigma_{j}\}_{j=1}^{r}$.
@@
$\operatorname{tr}(S_{\lambda})=\sum_{j=1}^{r}\frac{\sigma_{j}}{\sigma_{j}-\lambda},$
which is continuous and strictly increasing in $\lambda$. Moreover, $\operatorname{tr}(S_{0})=r$ and $\lim_{\lambda\uparrow\sigma_{\min}(\hat{\Sigma})}\operatorname{tr}(S_{\lambda})=+\infty$. Therefore, for any $\kappa>\tfrac{r}{n}$, the equation $\operatorname{tr}(S_{\lambda})=\kappa n$ has a unique solution.
@@
Using spectral decomposition and termwise differentiation, $\tfrac{d}{d\lambda}\operatorname{tr}(S_{\lambda})=\sum_{j}\tfrac{\sigma_{j}}{(\sigma_{j}-\lambda)^{2}}>0$, so monotonicity holds. Divergence at $\lambda\to\sigma_{\min}^{-}$ guarantees the intermediate value theorem applies. ∎
First-order Approximation for Small $\lambda$
@@
$\operatorname{tr}(S_{\lambda})=\sum_{j=1}^{r}\left(1+\frac{\lambda}{\sigma_{j}}+O(\lambda^{2})\right)=r+\lambda\sum_{j=1}^{r}\frac{1}{\sigma_{j}}+O(\lambda^{2}).$
Thus, an initial guess for $\lambda$ is given by
$\lambda^{(0)}\;\approx\;\frac{\kappa n-r}{\sum_{j=1}^{r}\tfrac{1}{\sigma_{j}}}\quad(\text{valid if }>0).$
Numerical Solution (Monotonic Equation)
@@
and update via $\lambda\leftarrow\lambda-\frac{g(\lambda)}{g^{\prime}(\lambda)}$, projecting back into $(-\infty,\,(1-\epsilon)\sigma_{\min}]$ to stay within the safe region.
Weighted Regularization $W\succeq 0$
For a general penalty $R(\theta)=\tfrac{1}{2}\theta^{\top}W\theta$, the trace becomes
$\operatorname{tr}(S_{\lambda})=\sum_{j}\frac{\sigma_{j}}{\sigma_{j}-\lambda w_{j}},$
so the same procedure applies with termwise substitutions $\sigma_{j}\leftarrow\sigma_{j},\ \lambda\leftarrow\lambda w_{j}$ (from a generalized eigenvalue perspective).
D.2 Power-law Decay Schedule
While degree-of-freedom targeting can increase $\lambda$ depending on the choice of $\kappa$ (especially when $r/n$ is small), we propose a more conservative alternative from a stability perspective: the decreasing power-law schedule:
$|\lambda(n)|=|\lambda(n_{0})|\left(\frac{n_{0}}{n}\right)^{\alpha},\qquad\alpha\geq 1.$
Proposition 23 (Recommendation of $\alpha=1$ via First-Order Balance).
@@ D.2 Power-law Decay Schedule
$\mathbb{E}[\mathcal{R}(\hat{f}_{\lambda})]=\mathbb{E}[\widehat{\mathcal{R}}_{S}(\hat{f}_{\lambda})]+\frac{2\tau^{2}}{n}\operatorname{tr}(S_{\lambda}),$
suppose for small $\lambda$ that $\operatorname{tr}(S_{\lambda})=r+\lambda S_{1}+O(\lambda^{2}),\quad S_{1}=\sum_{j}\tfrac{1}{\sigma_{j}},$ and that $\widehat{\mathcal{R}}_{S}(\hat{f}_{\lambda})=\widehat{\mathcal{R}}_{S}(\hat{f}_{0})-B_{1}\lambda+O(\lambda^{2})$ with $B_{1}>0$ representing the underfitting correction slope. Then the first-order term in test risk becomes
@@
Balancing the two terms in order of magnitude suggests that $\lambda(n)\propto n^{-1}$ is a natural choice.
@@
We compare the Taylor expansions of both terms and note that the variance penalty scales as $\tfrac{1}{n}$, so first-order balance yields $\lambda(n)=O(n^{-1})$. ∎
Remark 9 (Practical Guidelines).
The default $\alpha=1$ is recommended for general use. For noisier datasets or sharper spectra (many small $\sigma_{j}$), choosing $\alpha>1$ leads to faster decay, mitigating instability near the safety boundary. Degree-of-freedom targeting and power schedules are alternative strategies: the former targets constant complexity, while the latter emphasizes conservative stability.
@@ D.3 Numerical Stabilization Notes
Safety Margin and Clipping
@@
Initialization
Use the small-$\lambda$ approximation $\lambda^{(0)}\approx\tfrac{\kappa n-r}{\sum_{j}1/\sigma_{j}}$ as initial guess. If the value is negative, set $\lambda^{(0)}\leftarrow 0$. Spectral components can be approximated efficiently via randomized SVD or power iteration.
Line Search
When using gradient descent or quasi-Newton methods, apply external line search to ensure descent in $\hat{F}_{\lambda}(\theta)$. After each step, project $\lambda$ back into $(-\infty,(1-\epsilon)\sigma_{\min}]$ to maintain the safety region.
Newton Damping
In Newton updates $\lambda\leftarrow\lambda-\frac{g}{g^{\prime}}$, multiply by a damping factor $\eta_{\lambda}\in(0,1]$. If the new $\lambda$ exceeds the safe region, iteratively halve $\eta_{\lambda}$ until safety is restored.
Precision Management
When $(\sigma_{j}-\lambda)$ becomes very small, numerical precision loss becomes critical. In such cases, either (i) reduce the target $\kappa$ to push the root further away from the boundary, or (ii) cutoff small modes by excluding $\sigma_{j}<\delta$ from the trace approximation.
Log-Scale Tuning
For practical hyperparameter tuning, perform one-dimensional search in $\log_{10}\lambda$ space (e.g., golden section). Combine this with boundary clipping for robust and efficient convergence.
Appendix E Spectral Weighting and Noise Direction Control
E.1 Smoother under Weighted Regularization
Let the regularizer be defined as $R(\theta)=\tfrac{1}{2}\theta^{\top}W\theta$, where $W\succeq 0$. The empirical objective becomes
@@ E.1 Smoother under Weighted Regularization
and the normal equation is given by $(\hat{\Sigma}-\lambda W)\hat{\theta}_{\lambda}=\tfrac{1}{n}X^{\top}y$. Thus, the prediction $\hat{f}_{\lambda}=X\hat{\theta}_{\lambda}$ defines a linear smoother:
@@
Safety Condition
If $\hat{\Sigma}-\lambda W\succ 0$, then $\hat{F}_{\lambda}$ is strongly convex and $\hat{\theta}_{\lambda}$ is uniquely defined. Let $\rho_{\max}=\rho_{\max}(\hat{\Sigma},W)$ be the largest generalized eigenvalue of $\hat{\Sigma}v=\rho Wv$, then
$\lambda<\frac{1}{\rho_{\max}(\hat{\Sigma},W)}\quad\Longleftrightarrow\quad\hat{\Sigma}-\lambda W\succ 0.$
@@
Whitened Representation
When $\hat{\Sigma}\succ 0$, let $A:=\hat{\Sigma}^{-1/2}W\hat{\Sigma}^{-1/2}\succeq 0$, then:
$S_{\lambda}^{(W)}=\frac{1}{n}X\hat{\Sigma}^{-1/2}(I-\lambda A)^{-1}\hat{\Sigma}^{-1/2}X^{\top}=\frac{1}{n}Z(I-\lambda A)^{-1}Z^{\top},\quad Z:=X\hat{\Sigma}^{-1/2},$
with $Z^{\top}Z=nI$. This representation simplifies further derivations of degrees of freedom and variance.
E.2 Proof of Degrees of Freedom Formula
Proposition 24 (Degrees of Freedom under Weighted Regularization).
Under the above notation, let the eigenvalues of $A=\hat{\Sigma}^{-1/2}W\hat{\Sigma}^{-1/2}$ be $a_{1},\dots,a_{r}$ (where $r=\operatorname{rank}(X)$). Then, for $\lambda<1/\sigma_{\max}(A)$ (i.e., within the safety region),
$\operatorname{tr}(S_{\lambda}^{(W)})=\operatorname{tr}\big{(}(I-\lambda A)^{-1}\big{)}=\sum_{j=1}^{r}\frac{1}{1-\lambda a_{j}}.$
In particular, if $\hat{\Sigma}$ and $W$ are simultaneously diagonalizable, i.e., $\hat{\Sigma}=U\operatorname{diag}(\sigma_{j})U^{\top},\quad W=U\operatorname{diag}(w_{j})U^{\top},$ then
$\operatorname{tr}(S_{\lambda}^{(W)})=\sum_{j=1}^{r}\frac{\sigma_{j}}{\sigma_{j}-\lambda w_{j}}.$
@@ E.2 Proof of Degrees of Freedom Formula
$\operatorname{tr}(S_{\lambda}^{(W)})=\frac{1}{n}\operatorname{tr}\big{(}Z^{\top}Z(I-\lambda A)^{-1}\big{)}=\operatorname{tr}\big{(}(I-\lambda A)^{-1}\big{)}=\sum_{j}\frac{1}{1-\lambda a_{j}}.$
@@
A similar calculation yields:
$\operatorname{tr}\big{(}(S_{\lambda}^{(W)})^{2}\big{)}=\operatorname{tr}\big{(}(I-\lambda A)^{-2}\big{)}=\sum_{j}\frac{1}{(1-\lambda a_{j})^{2}}\quad\Big{(}\text{under simultaneous diagonalization: }\sum_{j}\frac{\sigma_{j}^{2}}{(\sigma_{j}-\lambda w_{j})^{2}}\Big{)},$
showing that variance inflation in low-denominator directions is directly governed by the choice of weights $w_{j}$.

### v2

@@ C.2 Proof of the Test Risk Improvement Theorem
Taking the difference between the two models gives
$\mathbb{E}\!\left[\mathcal{R}(\hat{f}_{\lambda})-\mathcal{R}(\hat{f}_{0})\right]=\mathbb{E}\!\left[\widehat{\mathcal{R}}_{S}(\hat{f}_{\lambda})-\widehat{\mathcal{R}}_{S}(\hat{f}_{0})\right]+\frac{2\tau^{2}}{n}\big{(}\operatorname{tr}(S_{\lambda})-\operatorname{tr}(S_{0})\big{)}.$
The second term equals $\tfrac{2\tau^{2}}{n}\Delta\operatorname{tr}$, while the first term equals $\Delta\widehat{\mathcal{R}}_{S}$. Under the assumption $\Delta\widehat{\mathcal{R}}_{S}>\tfrac{2\tau^{2}}{n}\Delta\operatorname{tr}$, the sum is negative, yielding the result. ∎
@@
At the equality threshold, the two risks coincide; below this threshold, test risk improvement occurs. The term $\tfrac{2\tau^{2}}{n}\operatorname{tr}(S_{\lambda})$ corresponds to a variance penalty, so if $\lambda$ drives the degrees of freedom excessively close to the safe boundary, this improvement effect vanishes.
@@ C.3 Monotonicity and Sensitivity Analysis
so $\operatorname{tr}(S_{\lambda})$ is monotone increasing in $\lambda$.
@@
Differentiate $\operatorname{tr}(S_{\lambda})=\sum_{j}\tfrac{\sigma_{j}}{\sigma_{j}-\lambda}$ term by term. ∎
Corollary 21 (Lipschitz-type upper and lower bounds).
@@
Moreover, for each term, $\frac{\sigma_{j}}{(\sigma_{j}-\lambda)^{2}}\geq\frac{s_{\min}}{(s_{\max}-\lambda)^{2}},$ so
@@
The sensitivity with respect to each $\sigma_{j}$ is
$\frac{\partial}{\partial\sigma_{j}}\Big{(}\frac{\sigma_{j}}{\sigma_{j}-\lambda}\Big{)}=-\frac{\lambda}{(\sigma_{j}-\lambda)^{2}}<0.$
Thus, larger small eigenvalues suppress the inflation of degrees of freedom, whereas smaller $\sigma_{j}$ make $\operatorname{tr}(S_{\lambda})$ more sensitive and prone to rapid growth.
@@
When many small $\sigma_{j}$ are present, the slope of $\operatorname{tr}(S_{\lambda})$ becomes steep and the variance penalty increases rapidly. It is therefore advisable to combine a safety margin $\lambda\leq(1-\epsilon)s_{\min}$ with a sample-size–dependent schedule $|\lambda(n)|\propto n^{-\alpha}$ so as to maintain the degrees of freedom at a controlled level per sample.
Appendix D Sample-size Scaling Rules (Regression)
D.1 Degrees-of-freedom Targeting Rule
This section establishes the existence and uniqueness of a solution $\lambda$ to the equation
@@ D.1 Degrees-of-freedom Targeting Rule
which aims to maintain the per-sample complexity at a target value $\kappa>0$. Here $S_{\lambda}=X(X^{\top}X-n\lambda I)^{-1}X^{\top}$, and $\{\sigma_{j}\}_{j=1}^{r}$ are the positive eigenvalues of $\hat{\Sigma}=\tfrac{1}{n}X^{\top}X$, with $r=\operatorname{rank}(X)$.
@@
$\operatorname{tr}(S_{\lambda})=\sum_{j=1}^{r}\frac{\sigma_{j}}{\sigma_{j}-\lambda}$
[…]
@@
From the spectral representation, term-wise differentiation gives $\tfrac{\partial}{\partial\lambda}\operatorname{tr}(S_{\lambda})=\sum_{j}\tfrac{\sigma_{j}}{(\sigma_{j}-\lambda)^{2}}>0$, establishing monotonicity. As $\lambda\to\sigma_{\min}^{-}$, the contribution of the smallest eigenvalue diverges. By continuity, monotonicity, and the unbounded limit, the intermediate value theorem ensures existence and uniqueness of the solution. ∎
First-order approximation for small $\lambda$ (initial guess).
@@
$\operatorname{tr}(S_{\lambda})=\sum_{j=1}^{r}\Big{(}1+\frac{\lambda}{\sigma_{j}}+O(\lambda^{2})\Big{)}=r+\lambda\sum_{j=1}^{r}\frac{1}{\sigma_{j}}+O(\lambda^{2}).$
Thus, an initial approximation for the target $\kappa n$ is
$\lambda^{(0)}\;\approx\;\frac{\kappa n-r}{\sum_{j=1}^{r}\tfrac{1}{\sigma_{j}}},\quad(\text{valid if positive}).$
Numerical solution (monotone equation).
@@
and update $\lambda$ by subtracting $g(\lambda)/g^{\prime}(\lambda)$. At each step, project $\lambda$ back into the safe region $(-\infty,(1-\epsilon)\sigma_{\min}]$ to preserve validity.
Weighted reward $W\succeq 0$.
If the reward is modified to $R(\theta)=\tfrac{1}{2}\theta^{\top}W\theta$, then
$\operatorname{tr}(S_{\lambda})=\sum_{j}\frac{\sigma_{j}}{\sigma_{j}-\lambda w_{j}}.$
Thus, the above procedure remains valid with the substitution $\lambda\mapsto\lambda w_{j}$ for each term, which corresponds to a generalized eigenvalue problem interpretation.
D.2 Derivation of the Power-law Schedule
Degrees-of-freedom targeting may result in $\lambda$ increasing with $n$, particularly when $r/n$ is small. As a conservative alternative from a stability perspective, we propose a decreasing power-law schedule:
$|\lambda(n)|=|\lambda(n_{0})|\Big{(}\frac{n_{0}}{n}\Big{)}^{\alpha},\qquad\alpha\geq 1.$
Proposition 24 (Recommendation of $\alpha=1$ from first-order balance).
@@ D.2 Derivation of the Power-law Schedule
for small $\lambda$,
$\operatorname{tr}(S_{\lambda})=r+\lambda S_{1}+O(\lambda^{2}),\qquad S_{1}=\sum_{j}\tfrac{1}{\sigma_{j}}.$
Also,
$\widehat{\mathcal{R}}_{S}(\hat{f}_{\lambda})=\widehat{\mathcal{R}}_{S}(\hat{f}_{0})-B_{1}\lambda+O(\lambda^{2}),$
where $B_{1}>0$ represents the slope correcting underfitting. The first-order term of the test risk difference is
@@
To balance these terms at the same order, $\lambda(n)\propto n^{-1}$ is the natural choice.
@@
Compare the Taylor expansion of the small-$\lambda$ terms with the $\tfrac{2\tau^{2}}{n}$ scaling. At the balance point, the effective order of $\lambda$ is $n^{-1}$. ∎
Remark 13 (Practical guideline).
Set $\alpha=1$ as the default. When the data are noisy or many small eigenvalues exist, yielding a sharp spectrum, choosing $\alpha>1$ for faster decay helps prevent blow-up near the boundary. Degrees-of-freedom targeting and power-law schedules serve as alternative rules: the former focuses on maintaining target complexity, while the latter emphasizes conservative stability.
@@ D.3 Numerical Stabilization Notes
Safety margin and clipping.
@@
Initialization.
Use the small-$\lambda$ approximation $\lambda^{(0)}\approx\tfrac{\kappa n-r}{\sum_{j}1/\sigma_{j}}$. If the approximation is negative, set $\lambda^{(0)}\leftarrow 0$. Spectra can be approximated efficiently via randomized SVD or the power method.
Line search.
When using gradient descent or quasi-Newton methods, employ an outer line search to guarantee descent of $\hat{F}_{\lambda}(\theta)$, and after each step project $\lambda$ back into $(-\infty,(1-\epsilon)\sigma_{\min}]$.
Newton damping.
For Newton updates, adjust $\lambda\leftarrow\lambda-g/g^{\prime}$ with a damping factor $\eta_{\lambda}\in(0,1]$. If the updated value exits the safe region, halve $\eta_{\lambda}$ iteratively until $\lambda$ re-enters the safe interval.
Precision management.
When $(\sigma_{j}-\lambda)$ becomes very small, numerical precision degrades. In such cases: (i) instead of decreasing $\epsilon$, reduce the target $\kappa$ to move the solution away from the boundary; or (ii) cutoff very small eigenmodes by excluding terms with $\sigma_{j}<\delta$ and approximate the degrees of freedom with the remaining components.
Log-scale tuning.
For practical hyperparameter tuning, performing a one-dimensional search in $\log_{10}\lambda$ space accelerates convergence, especially when combined with safety-margin clipping.
Appendix E Spectral Weighting and Control of Noise Directions
E.1 Linear smoother under weighted regularization
Let the reward be $R(\theta)=\tfrac{1}{2}\,\theta^{\top}W\theta$ with $W\succeq 0$. The empirical objective is
@@ E.1 Linear smoother under weighted regularization
and the normal equation is $(\hat{\Sigma}-\lambda W)\hat{\theta}_{\lambda}=\tfrac{1}{n}X^{\top}y$. Hence, the prediction $\hat{f}_{\lambda}=X\hat{\theta}_{\lambda}$ admits the linear smoother representation
@@
Safety condition.
If $\hat{\Sigma}-\lambda W\succ 0$, then $\hat{F}_{\lambda}$ is strongly convex and the minimizer $\hat{\theta}_{\lambda}$ is unique. In terms of the generalized eigenvalue problem $\hat{\Sigma}v=\rho\,Wv$ with maximum eigenvalue $\rho_{\max}$,
$\lambda<\frac{1}{\rho_{\max}\big{(}\hat{\Sigma},W\big{)}}\quad\Longleftrightarrow\quad\hat{\Sigma}-\lambda W\succ 0.$
@@
Whitened representation.
If $\widehat{\Sigma}\succ 0$, define $A:=\widehat{\Sigma}^{-1/2}W\widehat{\Sigma}^{-1/2}\succeq 0$ and $Z=X\widehat{\Sigma}^{-1/2}$. Then $Z^{\top}Z=nI$, and
$S^{(W)}_{\lambda}=\tfrac{1}{n}\,Z(I-\lambda A)^{-1}Z^{\top}.$
If $\widehat{\Sigma}$ is only positive semidefinite with rank deficiency, replace $\widehat{\Sigma}^{-1/2}$ with the Moore–Penrose pseudoinverse $\widehat{\Sigma}^{\dagger/2}$ on $\operatorname{col}(X)$, or alternatively add a small ridge $\varepsilon I$ and take the limit as $\varepsilon\to 0$.
E.2 Proof of the Degrees-of-freedom Formula
Proposition 25 (Degrees of freedom under weighted reward).
Let $A=\hat{\Sigma}^{-1/2}W\hat{\Sigma}^{-1/2}$ with eigenvalues $a_{1},\dots,a_{r}$ ($r=\operatorname{rank}(X)$). Then, in the safe region $\lambda<1/\sigma_{\max}(A)$,
$\operatorname{tr}\big{(}S_{\lambda}^{(W)}\big{)}=\operatorname{tr}\big{(}(I-\lambda A)^{-1}\big{)}=\sum_{j=1}^{r}\frac{1}{\,1-\lambda a_{j}\,}.$
In particular, if $\hat{\Sigma}$ and $W$ are simultaneously diagonalizable, i.e., $\hat{\Sigma}=U\operatorname{diag}(\sigma_{j})U^{\top},\,W=U\operatorname{diag}(w_{j})U^{\top},$ then
$\operatorname{tr}\big{(}S_{\lambda}^{(W)}\big{)}=\sum_{j=1}^{r}\frac{\sigma_{j}}{\sigma_{j}-\lambda w_{j}}.$
@@ E.2 Proof of the Degrees-of-freedom Formula
$\operatorname{tr}\big{(}S_{\lambda}^{(W)}\big{)}=\frac{1}{n}\operatorname{tr}\!\big{(}Z^{\top}Z\,(I-\lambda A)^{-1}\big{)}=\operatorname{tr}\!\big{(}(I-\lambda A)^{-1}\big{)}=\sum_{j}\frac{1}{1-\lambda a_{j}}.$
@@
A similar calculation gives
$\operatorname{tr}\big{(}(S_{\lambda}^{(W)})^{2}\big{)}=\operatorname{tr}\big{(}(I-\lambda A)^{-2}\big{)}=\sum_{j}\frac{1}{(1-\lambda a_{j})^{2}}\quad\Big{(}\text{or, under simultaneous diagonalization, }\sum_{j}\frac{\sigma_{j}^{2}}{(\sigma_{j}-\lambda w_{j})^{2}}\Big{)}.$
This shows that variance inflation in directions with small denominators is directly governed by the choice of weights $w_{j}$.